Chippawa Industries acquired a delivery truck on January 4, 2010. The total cost of the truck was $145,000. Chippawa estimated that the truck would be used for 8 years before being sold for an estimated $23,500. Chippawa uses the double-declining balance method of depreciation. The total depreciation expense for the year ended December 31, 2010 was:

18,125

15,188

30,375

36,250

Hunter Corporation beginning of year retained earnings balance was $28,300. The corporation declared and paid dividends of $19,400 during the year and ended the year with a $36,500 balance. The net income or loss for the year was:

(11,200)

8,200

27,600

17,100
Answer #1

Chippawa Industries

Under DDB method, salvage value is ignored

SLM rate = 1/ 8 = 12.5 %

DDB rate = 12.5 * 2 = 25 %

Dep, ( 2010 ) = 145000 * 25 % = $ 36250 ( Option D is correct )

Hunter Corporation

OPening RE + NI - Dividend = Closing RE

28300 + NI - 19400 = 36500

NI + 8900 = 36500

NI = $ 27600

OPtion C is correct.